Circular No. 5682

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVAE 1992bo AND 1992bp J. Maza, University of Chile (UC); and M. Hamuy, Cerro Tololo In-

teramerican Observatory, report the discovery of two apparent supernovae found on 20-min unfiltered IIa-O plates taken by G. Valladares with the CTIO Curtis Schmidt telescope. SN 1992bo was discovered by R. Antezana (UC), was at m_{pg} ∼ 17 on Dec. 20.06 UT, and is located 48′′ west and 55′′ south of the nucleus of the spiral galaxy ESO 352-G57 (α = 1^{h}19^{m}44^{s}.1, δ = −34^{o}27′29′′, equinox 1950.0). SN 1992bo was not visible (m_{pg} > 19.5) on a IIa-O plate taken on Nov. 30.1. SN 1992bp was discovered by M. Wischnjewsky (UC), was at m_{pg} ∼ 18 on Dec. 20.16, and is located ∼ 6′′ west and 1′′.5 south of the nucleus of a spiral galaxy at α = 3^{h}34^{m}22^{s}.3, δ = −18^{o}31′04′′. Confirmation of both supernovae was made by R. Aviles on CCD B, V, R, and I images obtained on Dec. 22.11 and 22.19 with the CTIO 0.9-m telescope.

SUPERNOVAE 1975V, 1976Q, 1977I, 1978L, 1990ak The Central Bureau has assigned letter designations to the following

apparent supernovae found on the SERC J and R surveys by R. H. Mc-Naught during comparison with U.K. Schmidt plates taken over the past two years. SN 1978L was on an R survey, all others on the J survey. The 1990 J plate is not part of the originally published survey. In no cases were other images found on additional UKST and ESO survey films. The positions are accurate to ∼ 0′′.4 in each coordinate (offsets measured from galaxies’ centers).

Object  Date (UT)  α_{1950}  δ_{1950}  Mag. Offsets
SN 1975V 1975 Aug. 29 20^{h}13^{m}13.^{s}99 −69^{o}12′58′′.5 19 5′′.7 W, 0′′.1 N
SN 1976Q 1976 Nov. 14 21 16 37.72 −67 26 08.6 18.5 0.8 W, 5.2 S
SN 1977I 1977 Aug. 18 1 05 42.75 −27 54 03.8 18 1.5 E, 20.4 S
SN 1978L 1978 Aug. 25 20 42 52.78 −53 30 10.3 18 7.0 W, 3.3 S
SN 1990ak 1990 May 1 13 22 06.10 −32 09 49.4 18 0.8 W, 10.8 S

SN 1975V is on the outer edge of what appears to be a spiral galaxy and does not appear on the ESO B survey (1975 June 13). SNe 1976Q and 1977I are within spiral arms; SN 1978L is on the outermost edge of ESO 412-G6 = MCG –05-03-026; SN 1990ak is in an elliptical galaxy.

1992 December 22 Daniel W. E. Green